OpenStudy (anonymous):

Help meeee! Please At a pumpkin patch, if Armando guesses the weight of his pumpkin within 0.3 ounces, he gets to take the pumpkin home for free. If his pumpkin weighs 4.9 pounds, which two equations can be used to find the minimum and maximum weights he can guess in order to get his pumpkin for free? a. x – 0.3 = 4.9 and x – 0.3 = –4.9 b. x + 0.3 = 4.9 and x + 0.3 = –4.9 c. x + 4.9 = 0.3 and x + 4.9 = –0.3 d. x – 4.9 = 0.3 and x – 4.9 = –0.3

OpenStudy (jack1):

4.9 = wt 4.9 + 0.3 = 5.2 4.9 - 0.3 = 4.6 so which 2 equations give the above answers for x?

OpenStudy (phi):

none of these work. they give us 4.9 pounds for the pumpkin and 0.3 ounces for the allowed error. If they meant 0.3 pounds, then one of the choices will be correct.

OpenStudy (anonymous):

What one should I guess then..? >.<

OpenStudy (jack1):

move to somewhere that uses metric system would be best choice... ;)

OpenStudy (anonymous):

Thanks @Jack1

OpenStudy (jack1):

i dont really get imperial sys @phi ... but guessing from the above that 1x ounce isnt 1 tenth of a pound?

OpenStudy (phi):

16 ounces to the pound. But the choices all imply 0.3 and 4.9 are in the same units. I would assume they meant 0.3 pounds to answer the question (but I would ask the teacher about this one)

OpenStudy (anonymous):

It's my online class xc

OpenStudy (jack1):

k k @Livie i think maths-wise, you're looking for equations you can rearrange to form x +0.3 = 4.9 and x - 0.3 = 4.9 but @phi is correct, whoever weote the question has about as much knowledge of imperial units as I do (so none)... so none of the answers are correct in the real sense

OpenStudy (anonymous):

Alrigt... But once I submit this assignment, and I guess it wrong, it'll give me a correct answer so I'll let you both know what it is.

OpenStudy (jack1):

hint: d

OpenStudy (anonymous):

It is D. That was the real answer...

OpenStudy (jack1):

a is -0,3 and - 0.3 b is +0,3 and +0.3 c gives -ve x values... has to be d

OpenStudy (anonymous):

Thanks for the help (: